# FlightBERT++: A Non-autoregressive Multi-Horizon Flight Trajectory Prediction Framework

### v2

@@ Introduction
Table 1 reports the overall performance of the proposed framework and comparison baselines. To investigate the robustness of these models with the horizons increase, the experimental results are divided into four (1, 3, 9, 15) different horizons, corresponding to 20 seconds, 1, 3, and 5 minutes trajectories in the future. It is demonstrated that the proposed FlightBERT++ framework achieves significant performance improvements against FlightBERT and outperforms other baselines in the MAE metric across all predicted attributes. For the RMSE metrics, thanks to the reduction of the BE representation bits, great RMSE reductions of the longitude, latitude, and altitude are obtained in the proposed framework. In addition, benefiting from the dedicated network design of the proposed framework, the error accumulation of the multi-horizon predictions is significantly decreased compared to competitive baselines.
@@
For the iterative multi-horizon prediction approaches, the KF-based model-driven approach suffers from a huge performance degradation with the increasing of prediction horizon due to the lack of real observations. Compared with KF-based approach, the data-driven based approaches achieved better performance across all prediction horizons. However, these approaches also suffer from error accumulation problems significantly in multi-horizon inference procedures. Benefiting from the capacity of the BE representation, the FlightBERT harvests comparable MAE performance in all prediction horizons. However, the RMSE of the FlightBERT is higher than other data-driven baselines due to the outliers caused by high-bit errors in the prediction. The LSTM+Attention achieves superior performance than iterative multi-horizon prediction approaches for long-term horizon predictions (in horizons 9 and 15). However, it also suffers from the performance degradation of the short-term horizons (horizon 1) in both the MAE and RMSE metrics.
@@
Computational Performance Evaluation
@@
FlightBERT++ 44.13 6.81
Table 2: Comparison of the computational performance.
The computational performance and the size of the model parameters are presented in Table 2. In this experiment, the prediction horizon is set to 15 for all models. As can be seen from the results, it is evident that the KF-based model achieves faster prediction speed due to its lower computational complexity and few parameters. Among the deep learning-based models, the direct multi-horizon approaches demonstrate substantial improvements in computational performance compared to the iterative multi-horizon approaches. In addition, it is worth noting that the model size of FlightBERT and FlightBERT++ is larger than comparison models. It can be attributed that the BE representations extend the dimension of the inputs and enable us to dedicatedly design the sophisticated neural architecture to capture the flight transition patterns. Moreover, the proposed FlightBERT++ harvests the fastest computational speed among deep learning based models, even with a tens larger model size than other baselines. This significant improvement in computational efficiency is primarily attributed to the design of the HACG module, enabling the FlightBERT++ to perform multi-horizon prediction in a non-autoregressive manner.
Visualization and Qualitative Analysis
@@
Figure 3: Visualization of the FTP results in selective flight scenarios, in which the altitude is measured in 10 $m$.
In this section, to better understand the learned flight transition patterns and qualitatively analyze the performance of different approaches, a total of 5 typical flight scenarios in the test set are selected to visualize the prediction results in Figure 3, including common flight scenarios (descending, climbing, and en-route) and complex flight patterns (turn, climbing and turn right). Each sample is visualized with 9 observation trajectory points (inputs), ground truth, and 15 predicted trajectory points generated by different methods.
As shown in Figure 3, the FlightBERT++ achieves superior performance over comparison baselines for both common flight scenarios and complex flight patterns. It is observed that the proposed model also exhibits the ability to estimate the flight intents in future horizons (Figure 3d, 3e). This indicates that the FlightBERT++ not only captures the flight dynamics from the observations but also learns typical flight patterns, such as fixed waypoints of turns or descents, from a substantial amount of historical trajectories in training samples. Moreover, the FlightBERT++ shows lower error accumulation during the multi-horizon prediction process, indicating significant reliability over comparison approaches. Considering the enhancement of downstream tasks, such as conflict detection and airspace planning, the FlightBERT++ can be regarded as a powerful tool to improve the overall efficiency and safety of ATC operations.
Insights
Ablation study
Exp. Models Hor. MAE $\downarrow$ MAPE (%) $\downarrow$ RMSE $\downarrow$ MDE $\downarrow$
@@
C1 FlightBERT++ (without ASA module) 1 0.0018 0.0018 1.18 0.0016 0.0067 0.21 0.0039 0.0119 12.10 0.32
3 0.0033 0.0032 2.28 0.0031 0.0119 0.42 0.0071 0.0138 12.56 0.57
9 0.0080 0.0077 5.25 0.0074 0.0285 0.94 0.0180 0.0243 17.89 1.34
15 0.0123 0.0118 7.16 0.0114 0.0439 1.31 0.0274 0.0339 22.38 2.05
C2 FlightBERT++ (without differential prompted mechanism) 1 0.0019 0.0018 1.27 0.0018 0.0071 0.23 0.0044 0.0116 12.18 0.34
3 0.0035 0.0034 2.56 0.0033 0.0126 0.46 0.0084 0.0138 12.97 0.60
9 0.0091 0.0086 6.71 0.0085 0.0318 1.15 0.0227 0.0267 20.56 1.51
15 0.0153 0.0144 10.66 0.0143 0.0534 1.81 0.0386 0.0409 28.59 2.53
Table 3: The experimental results of the ablation study.
In this section, to better understand the contributions of the designed components in the FlightBERT++ framework, an additional group C experiment is presented to conduct the ablation study. In experiment C1, a naive sum operator is applied to replace the ASA module in the trajectory encoder, while the differential prompt mechanism is removed from the DPD in experiment C2.
Figure 4: The histogram of the distance error.
@@
As can be seen in Table 3, all the designed network components make expected contributions to the FlightBERT++ framework. In experiment C1, the model suffers from more considerable error accumulation than the original FlightBERT++ with the increasing of prediction horizon. The weighted sum operation of the ASA module is believed to be more effective in capturing informative flight dynamics and temporal correlations from the observation trajectory sequence. Furthermore, in experiment C2, it can be found that the differential prompted mechanism is critical to the proposed FlightBERT++, especially in the long horizon prediction settings. It can be attributed that the differential prediction paradigm may lose some geographical and kinematical features of the flight trajectory, which makes it difficult to learn the differential transition patterns implicitly (without prompting) for the long-horizon prediction conditions. Therefore, the proposed differential prompted mechanism significantly contributes to the overall performance of the proposed FlightBERT++.
Error analysis (FlightBERT++ v.s. FlightBERT)
@@
To investigate the outliers caused by high-bit errors of BE representation, the distribution of the distance errors between the FlightBERT and FlightBERT++ is visualized by a histogram in Figure 4. It is evident that the FlightBERT framework still outputs larger distance error values in a certain number of samples. In contrast, thanks to the design of the differential prediction paradigm, the predictions of the FlightBERT++ show a significant reduction in outliers, which further supports our motivation to overcome the high-bit errors.
Conclusion

### v4

@@ V-A1 Overall Performance of FTP
Table I reports the overall performance of the proposed framework and comparison baselines. To investigate the robustness of these models with the horizons increase, the experimental results are divided into four (1, 3, 9, 15) different horizons, corresponding to 20 seconds, 1, 3, and 5 minutes trajectories in the future. It is demonstrated that the proposed FlightBERT++ framework achieves significant performance improvements against FlightBERT and outperforms other baselines in the MAE and MAPE metrics across all predicted attributes. Thanks to the proposed GC representation and the reduction of the bits by the differential prediction paradigm, the outliers caused by the high-bit misclassifications are effectively eliminated, leading to substantial reductions in RMSE for longitude, latitude, and altitude within the proposed framework. In addition, benefiting from the dedicated network design of the proposed framework, the error accumulation of the multi-horizon predictions is significantly decreased compared to competitive baselines. Furthermore, the MDE curve of the proposed FlightBERT++ and comparison baselines are illustrated in Figure 5. It can be found that the proposed FlightBERT++ shows the lowest MDE accumulation with the increase of the prediction horizon.
Notably, compared to the AAAI conference version, the FlightBERT++ in this paper achieves over 29.0% MDE reduction in 1 and 3 prediction horizons, as well as 24.8% and 21.8% MDE reduction in 9 and 15 prediction horizons. These experimental results further demonstrate the effectiveness of the proposed new GC representation in the MBC-based FTP task.
Figure 5: The comparison of the MDE curve over the 1 to 15 prediction horizons for different models.
@@
For the iterative multi-horizon prediction approaches, the KF-based model-driven approach suffers from a huge performance degradation with the increasing of the prediction horizon. The primary reason is that the KF-based FTP approach needs to update the parameters of the state equation dynamically to correct the prediction errors. However, the loss of real-time observation updates procedure leads to serious error accumulation in the multi-horizon prediction process. Compared to the KF-based approach, the data-driven based approaches achieved better performance across all prediction horizons. The Transformer-based model achieves the desired performance of the 1- and 3-horizon predictions but fails to output more accurate forecasting in the 9- and 15-horizon procedures. In contrast, the performance of the LSTM is slightly better than Transformer-based models in long-term horizons but loses the precision of the output in 1 horizon. It can be attributed that the Transformer-based models are more sensitive to prediction errors of pseudo-observations, especially in large error accumulation conditions, while the LSTM-based model makes it hard to fit complex flight patterns of the diverse flight phases for better performance. Benefiting from the capacity of the BE representation, the FlightBERT obtains better MAE performance in all prediction horizons. However, the RMSE of the FlightBERT is higher than other data-driven baselines due to the outliers caused by high-bit misclassification in the prediction. While the WTFTP demonstrates superior performance within 3 prediction horizons compared to all baselines, it suffers from heavy error accumulation during the iterative multi-horizon inference process.
The direct multi-horizon prediction baselines outperform iterative multi-horizon prediction approaches for long-term horizon predictions, particularly in horizons 9 and 15. However, the LSTM+Attention model suffers from the performance degradation of the short-term horizons (horizon 1) across both MAE and RMSE metrics. This result demonstrates that the LSTM+Attention approach primarily captures the global trajectory trends and dependencies over long-term horizons, but it falls short in effectively learning local and short-term fine-grained features. In contrast, the Transformer-Seq2Seq model achieves comparable results across both short and long prediction horizons, indicating its capability to effectively capture both global and local trajectory transition patterns.
@@
V-A2 Computational Performance Evaluation
TABLE II: Comparison of the computational performance.
@@ V-A2 Computational Performance Evaluation
Transformer-Seq2Seq 0.76 33.53
FlightBERT++ 29.96 6.95
The computational performance and the size of the model parameters are presented in Table II. To compare the prediction efficiency in multi-horizon settings, the number of the prediction horizon ($h$ in the Eq. (15)) is set to 15 for both FlightBERT++ and baseline models.
As can be seen from the results, it is evident that the KF-based model achieves faster prediction speed due to its lower computational complexity and few parameters. Among the deep learning-based models, the direct multi-horizon approaches demonstrate substantial improvements in computational performance compared to the iterative multi-horizon approaches. In addition, it is worth noting that the model size of FlightBERT and FlightBERT++ is larger than comparison models. It can be attributed that the BE representations extend the dimension of the inputs and enable us to dedicatedly design the sophisticated neural architecture to capture the flight transition patterns. Moreover, the proposed FlightBERT++ harvests the fastest computational speed among deep learning based models, even with a tens larger model size than other baselines. This significant improvement in computational efficiency is primarily attributed to the design of the HACG module, enabling the FlightBERT++ to perform multi-horizon prediction in a non-autoregressive manner. In summary, the experimental results demonstrate that FlightBERT++ exhibits higher computational efficiency, making it well-suited for supporting real-time forecasting in ATC environments.
Compared with the conference version [10], we streamline the network of the FlightBERT++ by setting the number of Transformer layers to 2 in the DPD, thereby reducing the number of parameters and computational complexity. However, it is worth noting that a slight increase in MTC is observed (6.95 ms v.s. 6.81 ms) in this version due to the additional encoding process of the GC representation.
V-B Visualization and Qualitative Analysis
In this section, to better understand the learned flight transition patterns and qualitatively analyze the performance of different approaches, a total of six typical flight scenarios in the test set are selected to visualize the prediction results. The visualization of selective samples is shown in Figure 6, including common flight scenarios (descending, climbing, and en-route) and complex flight patterns (turn, climbing and turn right, descending and maintain). Each sample is visualized in a 3D space with 9 observation trajectory points (inputs), ground truth, and 15 predicted trajectory points generated by the proposed FlightBERT++ framework and baseline models.
@@ V-B Visualization and Qualitative Analysis
(f) Descending and maintain
Figure 6: Visualization of the trajectory prediction results of selective flight scenarios, in which the altitude is measured in 10 $m$. Note that certain inaccurate predictions made by baselines are removed from figures to enhance readability.
As shown in Figure 6, the FlightBERT++ achieves superior performance over comparison baselines for both common flight scenarios and complex flight patterns. It is observed that the proposed model also exhibits the ability to estimate the flight intents in future horizons (Figure 6d, 6e, and 6f). This indicates that the FlightBERT++ not only captures the flight dynamics from the observations but also learns typical flight patterns, such as fixed waypoints of turns or descents, from a substantial amount of historical trajectories in training samples. Moreover, the FlightBERT++ shows lower error accumulation during the multi-horizon prediction process, indicating significant reliability over comparison approaches. Considering the enhancement of downstream tasks, such as conflict detection and airspace planning, the FlightBERT++ can be regarded as a powerful tool to improve the overall efficiency and safety of ATC operations.
For the LSTM+Attention and Transformer-Seq2Seq models, visualizations reveal that it is able to capture future flight trends for certain scenes. However, the LSTM+Attention model exhibits relatively larger prediction errors within 1 to 2 horizons (e.g., Figure 6b, 6c, and 6d), as that of the findings in the quantitative analysis. One possible reason is that the decoder of the LSTM+Attention model generates the 1 horizon predictions using a fixed $<\!\!sos\!\!>$ token, which brings a significant challenge in generating diverse and high-accurate results.
Among the iterative multi-horizon prediction approaches, it is observed that the error accumulation is enlarged with the increase of the prediction horizon. These models are able to predict the flight status accurately within 1 to 3 horizons but suffer from larger errors in longer horizons. Notably, the LSTM and Transformer models exhibit significant error accumulation, particularly in the altitude dimension (e.g., Figure 6c, 6d, and 6f). The Kalman filter-based method can achieve satisfactory prediction performance only in simple or transition patterns with linear variation flight scenarios (e.g., Figure 6a, 6b, and 6c). In addition, the FlightBERT model outperforms other iterative multi-horizon prediction approaches in most flight scenarios. However, the ability of flight trends perception is limited by the single-horizon prediction training strategy, making it challenging to precisely estimate future flight intents in complex or maneuvering flight phase (e.g., Figure 6d, 6e).
TABLE III: The experimental results of the ablation study.
Exp. Models Horizon MAE MAPE (%) RMSE MDE
@@
C1 FlightBERT++ (using BE representation) 1 0.0017 0.0017 1.14 0.0016 0.0062 0.22 0.0038 0.0074 9.06 0.31
3 0.0032 0.0031 2.32 0.0030 0.0115 0.43 0.0070 0.0122 10.91 0.55
9 0.0076 0.0074 5.60 0.0071 0.0273 1.04 0.0172 0.0228 17.65 1.29
15 0.0117 0.0112 8.04 0.0109 0.0417 1.51 0.0262 0.0319 23.70 1.96
C2 FlightBERT++ (w/o differential prompted mechanism) 1 0.0014 0.0015 1.10 0.0013 0.0056 0.19 0.0039 0.0115 12.07 0.26
3 0.0027 0.0026 2.23 0.0025 0.0099 0.39 0.0078 0.0131 12.57 0.46
9 0.0072 0.0069 6.15 0.0067 0.0256 1.04 0.0216 0.0244 19.73 1.20
15 0.0126 0.0118 9.96 0.0117 0.0439 1.65 0.0369 0.0378 27.57 2.07
C3 FlightBERT++ (w/o ASA module) 1 0.0013 0.0013 1.01 0.0012 0.0051 0.18 0.0032 0.0114 11.98 0.23
3 0.0024 0.0024 1.96 0.0022 0.0089 0.35 0.0060 0.0129 12.15 0.41
9 0.0060 0.0059 4.66 0.0056 0.0221 0.83 0.0160 0.0225 16.74 1.01
15 0.0095 0.0094 6.43 0.0088 0.0348 1.16 0.0248 0.0318 20.77 1.58
C4 FlightBERT++ (using Linear-based TPE) 1 0.0013 0.0013 0.99 0.0012 0.0050 0.17 0.0031 0.0113 11.96 0.23
3 0.0023 0.0023 1.94 0.0022 0.0088 0.53 0.0058 0.0126 12.10 0.40
9 0.0059 0.0058 4.41 0.0055 0.0217 0.83 0.0156 0.0216 16.63 0.99
15 0.0093 0.0093 6.45 0.0087 0.0344 1.18 0.0242 0.0306 20.70 1.56
FlightBERT++ (Proposed) 1 0.0012 0.0012 0.93 0.0011 0.0046 0.17 0.0031 0.0071 8.86 0.22
3 0.0023 0.0023 1.92 0.0022 0.0086 0.35 0.0059 0.0115 10.05 0.39
9 0.0058 0.0058 4.69 0.0054 0.0215 0.84 0.0156 0.0213 15.58 0.97
15 0.0091 0.0090 6.59 0.0085 0.0336 1.19 0.0239 0.0298 20.25 1.54
V-C Ablation study
The efficiency and effectiveness of the proposed FlightBERT++ framework are demonstrated in the above sections by the qualitative and quantitative evaluations. In this section, to confirm the contributions of the designed components in the FlightBERT++ framework, an additional group C experiment is introduced to conduct the ablation study. Specifically, the BE representation is applied to construct the FlightBERT++ framework to evaluate the effectiveness of the proposed GC representation in experiment C1. The differential prompt mechanism is removed from the DPD in experiment C2, while a naive sum operator is applied to replace the ASA module in the trajectory encoder in experiment C3. Furthermore, in experiment C4, to validate the effectiveness of the proposed Conv1D-based channel-mix TPE module, a set of linear layers [9] is employed to alternate the Conv1D layer to build a channel-independent TPE (Linear-based TPE) module in the FlightBERT++. In the above experiments, expected the validated modules, other experimental settings remain consistent with those outlined in Section IV to ensure experimental fairness and comparability.
@@ V-C Ablation study
The experimental results are presented in Table III, in which the w/o represents the FlightBERT++ implemented without specific modules. It can be found from the results that all the designed components make expected contributions to the FlightBERT++ framework. In experiment C1, it is observed that significant performance improvement is obtained by using the proposed GC representation compared to the BE representation. Furthermore, in experiment C2, it can be found that the differential prompted mechanism is critical to the proposed FlightBERT++, especially in the long-horizon prediction conditions. On the one hand, the differential prediction paradigm of the proposed FlightBERT++ may lose some geographical and kinematical features of the flight trajectory, which brings a challenge in learning spatial features of the output sequence. On the other hand, it is difficult to learn the differential transition patterns implicitly (without prompting) for the long-horizon prediction conditions. Therefore, the proposed differential prompted mechanism significantly contributes to the overall performance of the proposed FlightBERT++. In experiment C3, the model suffers from more considerable error accumulation than the original FlightBERT++ with the prediction horizon increasing. The weighted sum operation of the ASA module is believed to be more effective in capturing informative flight dynamics and temporal correlations from the observation trajectory sequence. As can be seen from the results of experiment C4, the model implemented by linear-based TPE is just slightly inferior to the original FlightBERT++ (Conv1D-based TPE), which also exhibits the robustness of the GC representations.
V-D Insights
V-D1 Representation ability analysis
(a) FlightBERT++
(b) LSTM+Attention
(c) Transformer-Seq2Seq
Figure 7: Visualization of the trajectory embeddings in 2D space via the TSNE tool.
To better understand the representation ability of the proposed FlightBERT++, a t-Distributed Stochastic Neighbor Embedding (TSNE) tool is employed to visualize the trajectory embeddings in a two-dimensional (2D) space. Specifically, we randomly selected 10 flight trajectories (T1 to T10 marked in Figure 7) in the test set. In succession, a total of 20 trajectory segments (containing 9 trajectory points for each segment) are extracted from each flight trajectory with a shift length of 1. The trajectory segments are further fed into the FTP models to extract the trajectory-level embeddings (e.g., $\mathbf{Traj}_{enc}$) and map these embeddings into 2D space by the TSNE tool.
The visualization of the trajectory embeddings extracted by the FlightBERT++, LSTM+Attention and Transformer-Seq2Seq models are presented in Figure 7a, 7b and 7c, respectively. In Figure 7, each colored point denotes a certain trajectory embedding, while different trajectories are depicted by individual colors. It is clear that the FlightBERT++ can capture discriminative features for different flight trajectories. Impressively, the proposed FlightBERT++ learns strong discriminative temporal features among different trajectory segments (even with minor changes of 1 shift length), as well as distinguishes the trajectory embeddings clearly in embedding space among different trajectories.
In contrast, the LSTM+Attention and the Transformer-Seq2Seq models only distinguish the embeddings of the different trajectories but hard to provide the required temporal discriminability among similar trajectory segments in the embedding space. In summary, the visualization clearly illustrates that the proposed FlightBERT++ has a stronger representation ability of trajectory sequence, which also provides interpretability for the superior performance of the proposed approach.
V-D2 Error analysis (FlightBERT++ v.s. FlightBERT)
Figure 8: An example of the outliers in FlightBERT caused by high-bit misclassification. The outliers are marked using the red circle.
To intuitively illustrate the outliers caused by high-bit misclassification in the BE representation, a case study of the multi-horizon prediction by FlightBERT and FlightBERT++ is visualized in Figure 8. Obviously, the high-bit misclassifications occurred in the longitudinal dimension during the $3^{rd}$ and $10^{th}$ prediction horizons, leading to huge prediction errors in the following horizons. In contrast, the FlightBERT++ performs the prediction more precisely by incorporating the proposed non-autoregressive multi-horizon forecasting mechanism, differential prediction, and GC representations. To further investigate the predicted outliers by FlightBERT and FlightBERT++, the absolute errors of LLA attributes and deviation errors are presented by a set of boxplots in Figure 9. The prediction errors of $1^{st},3^{rd},9^{th},15^{th}$ output by the FlightBERT and FlightBERT++ are illustrated in the boxplots, respectively. In order to present the box and whisker of each prediction horizon more clearly, the outliers are removed from the Figures due to the larger number of samples in the test set.
Figure 9: The boxplots of the absolute error and deviation error for FlightBERT and FlightBERT++.
@@ V-D1 Representation ability analysis
As can be seen from the boxplots, compared to the FlightBERT, FlightBERT++ achieves significant performance improvements, especially in longer prediction horizons. By analyzing the experimental results, it is found that the FlightBERT model outputs a considerable number of outliers, leading to a higher RMSE. To validate this, the distribution of the deviation errors between the FlightBERT and FlightBERT++ is visualized by a histogram in Figure 10. It is evident that the FlightBERT framework still outputs larger deviation error values in a certain number of samples. In contrast, thanks to the design of the GC representation and differential prediction paradigm, the predictions of the FlightBERT++ show a significant reduction in outliers, which further supports our motivation to address the high-bit misclassifications of the BE representation.
Figure 10: The comparison of the deviation error between FlightBERT and FlightBERT++ via histogram.
V-D3 Evaluation in different flight phases
To explore the robustness of the proposed FlightBERT++ in different flight phases, the trajectory segments in the test set are further categorized into three subsets according to the flight trends, including climbing, descending/approaching, and en-route. In general, the flight transition patterns in the en-route phase are close to the linear, while the climbing and descending/approaching phases are considered as maneuvering operations.
The experimental results are illustrated in Table IV. As expected, the FlightBERT++ achieves superior performance in the en-route phase, particularly in the attribute of altitude, which benefits from the linear motion patterns in this phase. It can be primarily attributed to two key factors: i) The transition patterns in the en-route phase are relatively linear and easy to capture, resulting in accurate predictions. ii) The training set comprises a large proportion of en-route samples, as most flights spend a considerable duration in the en-route phase during the entire flight procedure. It is also observed that the performance of the model suffers from a slight reduction in the climbing and descending phase due to the maneuvering operations. In summary, the proposed FlightBERT++ demonstrates the ability to effectively capture different flight dynamics and perform the FTP task across various flight phases, allowing it a promising and versatile approach in ATC- and FTP-related applications.
TABLE IV: The experimental results evaluated in different flight phases of proposed FlightBERT++.
Flight Phase Horizon MAE MAPE (%) RMSE MDE
Lon Lat Alt Lon Lat Alt Lon Lat Alt
Climbing 1 0.0018 0.0020 4.68 0.0017 0.0072 1.10 0.0031 0.0036 6.59 0.34
3 0.0034 0.0042 8.22 0.0032 0.0155 1.76 0.0060 0.0081 11.76 0.66
9 0.0095 0.0118 16.43 0.0090 0.0435 2.96 0.0178 0.0224 23.67 1.81
15 0.0152 0.0178 23.89 0.0143 0.0656 3.85 0.0291 0.0319 34.42 2.80
En-route 1 0.0011 0.0011 0.03 0.0011 0.0043 0.0033 0.0049 0.0124 0.33 0.20
3 0.0020 0.0019 0.07 0.0018 0.0074 0.0075 0.0071 0.0138 0.83 0.33
9 0.0044 0.0043 0.20 0.0041 0.0165 0.0231 0.0131 0.0197 2.90 0.73
15 0.0068 0.0066 0.32 0.0063 0.0251 0.0341 0.0185 0.0258 5.84 1.11
Descending Approaching 1 0.0012 0.0014 2.08 0.0011 0.0049 0.42 0.0027 0.0024 3.51 0.23
3 0.0026 0.0028 4.32 0.0025 0.0098 0.90 0.0074 0.0055 7.73 0.46
9 0.0089 0.0083 12.14 0.0085 0.0290 2.67 0.0248 0.0189 20.34 1.43
15 0.0149 0.0135 18.23 0.0142 0.0470 4.43 0.0387 0.0271 27.94 2.33
VI Conclusion and Future Work